Tohban Report 2013-02-27

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 20 Feb 2013
End Date: 27 Feb 2013
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity was low-moderate this week with only 10 C class flares. There are currently 7 active regions on the disk, although 4 of these will rotate over the limb within the next 36 hours. STEREO B shows a couple of active regions which will rotate onto the disk over the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     20    10     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     14     5     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      5     3     0     0

There were RHESSI flares/GOES flares 206 / 30 over the time range 20-Feb-13 27-Feb-13

Memory Management

The SSR is emptying at the end of the daily pass set, with a maximum of about 25%.

Cold plate 1 and 2 temperatures are at 111 K and 110 K respectively.

The cryocooler is nominal.

Spacecraft Status

Data Gaps

There were no data gaps this week.

Detector issues

The front slow threshold on detector 2 was lowered on days 052 and 053 (21st and 22nd February).

The front slow valid counts were around 1200 at 0x38, a little higher than the typical ~800 seen during previous orbits. However during recent passes the counts are around ~600. The variation may be caused by orbital variation or as a result of the pass occurring while the light leak in detector 2 is present. The new tohban will keep an eye on this and will observe the situation for another week.

Note we started at 0x50 on day 052.

2013-052-18:00:24 /IDPULOAD VALUE=0x4C

2013-052-18:02:13 /IDPULOAD VALUE=0x48

2013-052-18:06:15 /IDPULOAD VALUE=0x44

2013-053-16:10:30 /IDPULOAD VALUE=0x40

2013-053-16:11:20 /IDPULOAD VALUE=0x3C

2013-053-16:12:52 /IDPULOAD VALUE=0x38

2013-053-17:51:28 /IDPULOAD VALUE=0x34

2013-053-17:52:53 /IDPULOAD VALUE=0x36

2013-053-17:56:19 /IDPULOAD VALUE=0x38

The detector remains unsegmented.

Other notes

A spin up was carried out on the 26th February between ~17:00 UT and ~20:00UT.

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? No problems. The front slow threshold on detector 2 was lowered (see above).
Personal tools